# Modeling Binary Lenses and Sources with the BAGLE Python Package

### v1

@@ Abstract
Gravitational microlensing is a powerful tool that can be used to find and measure the mass of isolated and dark compact objects. In many microlensing events, the lens, the source, or both may be a binary system. Therefore, in this study we present lensing equations for binary source and lens models in the Bayesian Analysis of Gravitational Lensing Events (BAGLE) Python microlensing package. The new binary source and lens models in BAGLE account for the complete Keplerian orbit. BAGLE also includes binary models that approximate the orbital motion as linear or accelerating motion of the secondary companion; these are useful when the orbit has a very low eccentricity or the orbital period is much longer than the microlensing timescale. The model parameterizations based on these binary lensing equations will enable joint fitting of photometric and astrometric data sets. Consequently, binary microlensing events with complex astrometric trajectories can be used to break several microlensing degeneracies that plague photometry-only microlensing modeling. These binary models will be used to fit microlensing event data from the Vera C. Rubin Observatory, the Nancy Grace Roman Telescope, and other surveys.
@@ 1 Introduction
Microlensing is sensitive to both close and wide separation binaries, unlike radial velocity and transit probes, which are more sensitive to closely separated binaries or planet+star systems. Recent simulations of Milky Way microlensing surveys show that 55% of observed microlensing events involve a binary star system (N. S. Abrams et al., 2025), so robust binary microlensing models are even more important than previously thought. Some microlensing events have timescales greater than the orbital period of the involved binary system. In these cases, photometric lightcurves and astrometric trajectories can only be properly modeled by accounting for the system’s orbital dynamics.
@@
There are numerous software packages that encode the math necessary to model and fit a microlensing event such as BAGLE (J. Lu et al., submitted), pyLIMA (E. Bachelet et al., 2017), VBMicrolensing (V. Bozza, 2010; V. Bozza et al., 2018, 2021, 2024), RTModel (V. Bozza, 2024) and MulensModel (R. Poleski & J. C. Yee, 2019). These packages are all publicly available and have their own strengths and weaknesses in terms of the types of event geometries they support, whether they include both photometry and astrometry in their models, and how accurate and efficient their model-fitting capabilities are. A detailed comparison of these packages for point-source, point-lens models is presented in (J. Lu et al., submitted). PyLIMA, VBMicrolensing and MulensModel all support binary lenses, binary sources and both (with orbital motion). Only RTModel and BAGLE support joint-fitting of photometric and astrometric datasets, although RTModel does not yet support joint-fitting for binary-source, binary-lens events.
@@
In this work, we introduce models for binary systems into the Bayesian Analysis of Gravitational Lensing Events (BAGLE) Python package [^1] https://github.com/MovingUniverseLab/BAGLE_Microlensing. This paper is a companion to (J. Lu et al., submitted), which introduces BAGLE and presents models for single-lens object and single-source star events, including point-source, point-lens (PSPL); and finite-source, point-lens (FSPL) models. Here we describe the BAGLE implementation of models for point-source, binary-lens (PSBL); binary-source, point-lens (BSPL); and binary-source, binary-lens (BSBL) systems. The different binary geometries of microlensing systems are introduced in §2. The general mathematical framework for modeling binary orbital motion in BAGLE is described in §3. The complete equations of motion are presented for BSPL (§4), PSBL (§5), and BSBL (§6). Each of these sections contains sub-sections where we present models for static binaries with fixed primary + secondary positions, secondary companions that move with linear or accelerating motions with respect to their primary, and binary systems with full Keplerian orbital motion. Model validation and comparison with other packages is presented in §7. In §8, we present example magnification maps and centroid shift maps from various BAGLE binary models. Changes to photometric lightcurves and astrometric trajectories with binary mass ratio, separation, and orbital parametres are also discussed in this section. Conclusions are presented in §9 along with planned BAGLE upgrades.
2 Binary Lens Geometries
@@
3 Binary Orbital Motion in BAGLE
Figure 2: Trajectory of a binary orbit at $\textit{i}=0\degree$ (binary disk is face-on) simulated using BAGLE for stationary proper motion. The orbit has an eccentricity of 0.8. The primary and secondary objects at the time of periastron passage have been highlighted.
@@ 3 Binary Orbital Motion in BAGLE
  • $\omega_{pri}$: The argument of periastron of the primary object’s orbit in degrees. The secondary companion is placed $180\degree$ across the primary’s argument of periastron.
@@
In BAGLE, the eight Keplerian elements presented above are used to estimate the Thiele-Innes constants as described in S. C. Koren et al. (2016), T. N. Thiele (1883) and presented, in detail, in Appendix A. Using these constants, the positions of the primary and the secondary companion over time are computed. Let $\boldsymbol{X(t)}_{\boldsymbol{p},\boldsymbol{com},\sun}$ be the primary object’s projected trajectory (on the sky plane) around the center of mass, with ${X(t)}_{{p,{com},E},\sun}$ and ${X(t)}_{{p,{com},N},\sun}$ being the East and North components, respectively. Let $\boldsymbol{X(t)}_{\boldsymbol{s},\boldsymbol{com},\sun}$ be the secondary companion’s trajectory around the center of mass with ${X(t)}_{{s,{com},E},\sun}$ and ${X(t)}_{{s,{com},N},\sun}$ representing the East and North components, respectively. Then, the following set of equations holds:
@@
Using Eqn 3, BAGLE can model circular and elliptical orbital trajectories around the center of mass. An instance of a Keplerian elliptical orbital trajectory (eccentricity of 0.8) with stationary proper motion is shown in Figure 2. Note that the equations in this section describe the motion around the center of mass. The proper motion of the center of mass in the observer’s frame of reference is accounted for separately.
4 Binary Sources and Point Lenses (BSPL)
In this section, we will discuss microlensing models for binary sources. We begin by discussing simple, static approximations in §4.1. Then, we expand to a discussion of linear and accelerated orbital approximations in §4.2. Finally, the full Keplerian solutions are presented in §4.3, followed by a discussion of microlensing equations in §4.4.
4.1 Static Approximation
In binary-source models, the components of the binary (i.e., the primary source and the secondary source companion) are initially at rest relative to each other. On the plane of the sky, in the Solar-System Barycenter (SSB) frame, at $t_{0,\sun}$, the binary source system initially has a fixed angular separation of:
@@ 4.1 Static Approximation
where $\boldsymbol{s_{S}}(t_{0,\sun})$ is the separation vector at $t_{0,\sun}$, $\boldsymbol{X}_{\boldsymbol{S_{s}},\sun}(t_{0,\sun})$ is the initial secondary source position on the sky at $t_{0,\sun}$, and $\boldsymbol{X}_{\boldsymbol{S_{p}},\sun}(t_{0,\sun})$ is the initial primary source position on the sky at $t_{0,\sun}$. Note, we use the Sun symbol, $\odot$, to indicate SSB coordinates. In the binary-source models, $t_{0,\sun}=t_{prim,0,\sun}$, i.e., $t_{0,\sun}$ is the time of closest approach between the primary source and the lens ($t_{prim,0,\sun}$). The angular separation between the primary and the secondary companion is fixed at all times. In other words, the proper motion of the secondary companion is assumed to be the same as the proper motion of the primary source.
Figure 3: Source and lens trajectories for linear (Top) and accelerated (Bottom) approximations of orbital motion in binary sources. We present the unlensed sources (solid lines) and the lensed images (dashed + dotted lines). For each source, there is a major and a minor image. The minor image is seen around the lens, and the major image is seen around the source. The green line is the flux-weighted average of lensed source positions, as observed on the sky. Note that $mag_{S,pri}$ = 16 and $mag_{S,sec}$ = 17. In both panels, the primary’s proper source motion $\boldsymbol{\mu}_{\boldsymbol{S},\sun}$ is [6 mas yr-1, 3 mas yr-1]; the secondary’s proper source motion relative to the primary source $\boldsymbol{\mu}_{\boldsymbol{S_{s}},\sun}$ is [9 mas yr-1, 7 mas yr-1]. The acceleration of the secondary source $\mathbb{a}_{\boldsymbol{S\mathrm{rel}},\sun}$ is [0.5 mas yr-2, -2 mas yr-2] in the lower panel. In both cases, the Einstein time ($t_{E,\sun}$)=269 days and $u_{0,\sun}$ =1.01. The lens mass is $10M_{\odot}$, and it is held stationary.
4.2 Linear and Accelerated Orbital Approximations
@@
where $\boldsymbol{\mu}_{\boldsymbol{S},\sun}$ is the proper motion of the primary source, and $\boldsymbol{\mu}_{\boldsymbol{S_{s}},\sun}$ is the proper motion of the secondary source calculated as $\boldsymbol{\mu}_{\boldsymbol{S_{s}},\sun}=\boldsymbol{\mu}_{\boldsymbol{S},\sun}+\boldsymbol{\Delta\mu}_{\boldsymbol{S_{s}},\sun}$. Furthermore, we account for the parallactic motion at time $t$ in the direction of the source-lens system $(\alpha,\delta)$, given by the difference of the Earth and Sun’s position, normalized by 1 AU. The parallactic motion is $\pi_{S}\boldsymbol{P}(t,\alpha,\delta)$ where $\pi_{S}$ is the maximum parallax amplitude $1/d_{S}$ and $\boldsymbol{P}(t,\alpha,\delta)$ is the actual parallax direction and fractional amplitude on the sky.
@@
An example of the linear and accelerated orbital approximations is presented in Figure 3, showing the change in Right Ascension ($\Delta\alpha^{*}$) and Declination ($\Delta\delta$) of the primary source (lensed and unlensed), secondary source (lensed and unlensed), and the lens. The lensed images are further categorized into the major (“+”) image and minor (“-”) image. These figures can be generated using BAGLE with the code lines below. For this example, we use BSPL_PhotAstrom_noPar_LinOrbs_Param2, which is a model for linear orbits. One can follow the same steps below to generate plots for accelerated orbits:
@@
3#Create a linear orbital model with photometric and astrometric parameters
4
5bsplorbits = model.BSPL_PhotAstrom_noPar_LinOrbs_Param2(
@@
15
16# Get resolved astrometry for unlensed source positions
17t_obs = np.arange(t0 - 10*tE, t0 + 10*tE, 1)
18bsplorbits.get_resolved_source_astrometry_unlensed(t)
19srce_pos_primary =xS_unlensed[:, 0, :]
20srce_pos_secondary = xS_unlensed[:, 1, :]
21
22# Get lens astrometry
23lens = bsplorbits.get_lens_astrometry(t)
24
25# Get unresolved astrometry for lensed source positions
26xS_lensed = bsplorbits.get_astrometry_shift(t)
27lensed_pos_pri = xS_lensed[:, 0, :]
28lensed_pos_sec = xS_lensed[:, 1, :]
@@
4.3 Full Keplerian Solutions
Figure 4: Source and lens trajectories for a simulated binary source point lens microlensing event involving Keplerian orbits. We present the unlensed sources (solid lines) and the lensed images (dashed + dotted lines). The green line is the flux-weighted average of lensed source positions, as observed on the sky. Note that $mag_{S,pri}$ = 16 and $mag_{S,sec}$ = 17. This event has the following Keplerian elements: $\omega_{pri}=30\degree$, $\Omega_{sec}=10\degree$, $\textit{i}=90\degree$, $e=0.6$, $\textit{P}=1000$ days, $\aleph_{pri}=3$ mas and $\aleph_{sec}=8$ mas. We simulate this event over $t_{E}=208.47$ days. The lens (8 $M_{\odot}$) is held stationary.
All binary-source models with Keplerian motion in BAGLE input the following Keplerian orbital parameters: $\omega_{pri}$, i, $\Omega_{sec}$, e (for elliptical orbits only), P, $t_{p}$, $\aleph_{pri}$, $\aleph_{sec}$. These models input an additional quantity $\boldsymbol{\mu}_{\boldsymbol{S,com},\sun}$, which is the proper motion of the source system’s center of mass in the observer’s frame of reference. $\boldsymbol{X}_{\boldsymbol{S_{com},0},\sun}$ is the position of the binary source’s center of mass at $t_{0,\sun}=t_{prim,0,\sun}$. While BAGLE inputs $t_{0,\sun}=t_{prim,0,\sun}$ as the time of closest approach between the primary source and the lens, and $u_{0,\sun}$ as the closest approach between the primary source and the lens, it can convert between different coordinate systems (e.g., to use the closest approach between the source system center of mass and the lens). Conversions for $t_{0,\sun}$ and $u_{0,\sun}$ are presented in Appendix B and C.
In Figure 2, we demonstrated the motion of the primary and its secondary companion around the center of mass at rest. After accounting for the proper motion of the center of mass, the primary and secondary source trajectories (including parallactic motion) are given by:

### v3

@@ Abstract
Gravitational microlensing is a powerful tool that can be used to find and measure the mass of isolated and dark compact objects. In many microlensing events, the lens, the source, or both may be a binary system. In this work, we introduce binary source and lens models into the gravitational lensing formalism encoded in the Bayesian Analysis of Gravitational Lensing Events (BAGLE) Python software package. These new binary models in BAGLE account for Keplerian orbits. We also add binary models with fewer parameters that describe the binary orbital motion as acceleration, linear, or stationary motion of the secondary companion; these are useful when the orbit has a very low eccentricity or the orbital period is much longer than the microlensing timescale. The model parameterizations based on these binary lensing equations enable joint-fitting of photometric and astrometric datasets. These binary models will be used to fit microlensing event data from the Vera C. Rubin Observatory, the Nancy Grace Roman Telescope, and other surveys.
@@ I Introduction
Microlensing is sensitive to both close and wide separation binaries, unlike radial velocity and transit probes, which are more sensitive to closely separated binaries or planet+star systems. Recent simulations of Milky Way microlensing surveys show that 55% of observed microlensing events involve a binary star system (N. S. Abrams et al., 2025) most of which can masquerade as single source-single lens events. This is significantly more than the $\sim 5-11\%$ of microlensing events with obvious binaries which are commonly reported (e.g. C. Alcock et al., 2000; M. Jaroszynski, 2002; P. Mróz et al., 2019), so robust binary microlensing models are even more important than previously thought. Some microlensing events have timescales greater than the orbital period of the involved binary system. In these cases, photometric lightcurves and astrometric trajectories can only be properly modeled by accounting for the system’s orbital dynamics (Dominik:1998a; Skowron_2011; Penny_2011; V. Bozza, 2010).
@@
There are numerous software packages that encode the math necessary to model and fit a microlensing event such as BAGLE (J. Lu et al., submitted), pyLIMA (E. Bachelet et al., 2017), VBMicrolensing (V. Bozza, 2010; V. Bozza et al., 2018, 2021, 2024), RTModel (V. Bozza, 2024) and MulensModel (R. Poleski & J. C. Yee, 2019). These packages are all publicly available and have their own strengths and weaknesses in terms of the types of event geometries they support, whether they include both photometry and astrometry in their models, and how accurate and efficient their model-fitting capabilities are. A detailed comparison of these packages for point-source, point-lens models is presented in (J. Lu et al., submitted). PyLIMA, VBMicrolensing and MulensModel all support binary lenses, binary sources and both (with orbital motion). Only RTModel (which relies on using VBMicrolensing) and BAGLE support joint-fitting of photometric and astrometric datasets, although RTModel does not yet support joint-fitting for binary-source, binary-lens events.
@@
In this work, we introduce models for binary systems into the Bayesian Analysis of Gravitational Lensing Events (BAGLE) Python package [^1] https://github.com/MovingUniverseLab/BAGLE_Microlensing. This paper is a companion to J. Lu et al. (submitted), which introduces BAGLE and presents models for single-lens object and single-source star events, including point-source, point-lens (PSPL); and finite-source, point-lens (FSPL) models. Here we describe the BAGLE implementation of models for point-source, binary-lens (PSBL); binary-source, point-lens (BSPL); and binary-source, binary-lens (BSBL) systems. The different binary geometries of microlensing systems are introduced in §II. The general mathematical framework for modeling binary orbital motion in BAGLE is described in §III. The complete equations of motion are presented for BSPL (§IV), PSBL (§V), and BSBL (§VI). Each of these sections contains sub-sections where we present models for static binaries with fixed primary + secondary positions, secondary companions that move with linear or accelerating motions with respect to their primary, and binary systems with full Keplerian orbital motion. Model validation and comparison with other packages is presented in §VII. In §VIII, we present example magnification maps and centroid shift maps from various BAGLE binary models. Changes to photometric lightcurves and astrometric trajectories with binary mass ratio, separation, and orbital parametres are also discussed in this section. Conclusions are presented in §IX along with planned BAGLE upgrades.
II Binary Lens Geometries
@@
III Binary Orbital Motion in BAGLE
Figure 2: Trajectory of a binary orbit at $\textit{i}=0\degree$ (orbital plane
is face-on) simulated using BAGLE for stationary proper motion. The orbit has an eccentricity of 0.8. The primary and secondary objects at the time of periastron passage have been highlighted.
@@ III Binary Orbital Motion in BAGLE
  • $\omega_{pri}$: The argument of periastron of the primary object’s orbit in degrees. The secondary companion is placed $180\degree$ across the primary’s argument of periastron. When $\textit{e}=0$, we set $\omega_{pri}=0$ since there is no periapsis.
@@
In BAGLE, the eight Keplerian elements presented above are used to estimate the Thiele-Innes constants as described in S. C. Koren et al. (2016), T. N. Thiele (1883) and presented, in detail, in Appendix B. Using these constants, the positions of the primary and the secondary companion over time are computed. Let $\boldsymbol{X(t)}_{\boldsymbol{p},\boldsymbol{com},\sun}$ be the primary object’s projected trajectory (on the sky plane) around the center of mass, with ${X(t)}_{{p,{com},E},\sun}$ and ${X(t)}_{{p,{com},N},\sun}$ being the East and North components, respectively. Let $\boldsymbol{X(t)}_{\boldsymbol{s},\boldsymbol{com},\sun}$ be the secondary companion’s trajectory around the center of mass with ${X(t)}_{{s,{com},E},\sun}$ and ${X(t)}_{{s,{com},N},\sun}$ representing the East and North components, respectively. Then, the following set of equations holds:
@@
Using Eqn III, BAGLE can model circular and elliptical orbital trajectories around the center of mass. An instance of a Keplerian elliptical orbital trajectory (eccentricity of 0.8) with stationary proper motion is shown in Figure 2. Note that the equations in this section describe the motion around the center of mass. The proper motion of the center of mass in the observer’s frame of reference is accounted for separately.
IV Binary Sources and Point Lenses (BSPL)
In this section, we will discuss microlensing models for binary sources. We begin by discussing simple, static approximations in §IV.1. Then, we expand to a discussion of linear and accelerated orbital approximations in §IV.2. Finally, the full Keplerian solutions are presented in §IV.3, followed by a discussion of microlensing equations in §IV.4.
IV.1 Static Approximation
In binary source models, the components of the binary (i.e., the primary source and the secondary source companion) are initially at rest relative to each other. On the plane of the sky, in the Solar-System Barycenter (SSB) frame, at $t_{0,\sun}$, the binary source system initially has a fixed angular separation of:
@@ IV.1 Static Approximation
where $\boldsymbol{s_{S}}(t_{0,\sun})$ is the separation vector at $t_{0,\sun}$, $\boldsymbol{X}_{\boldsymbol{S_{s}},\sun}(t_{0,\sun})$ is the initial secondary source position on the sky at $t_{0,\sun}$, and $\boldsymbol{X}_{\boldsymbol{S_{p}},\sun}(t_{0,\sun})$ is the initial primary source position on the sky at $t_{0,\sun}$. Note, we use the Sun symbol, $\odot$, to indicate SSB coordinates. In the binary source models, $t_{0,\sun}=t_{prim,0,\sun}$, i.e., $t_{0,\sun}$ is the time of closest approach between the primary source and the lens ($t_{prim,0,\sun}$). The angular separation between the primary and the secondary companion is fixed at all times.
Figure 3: Source and lens trajectories for linear (Top) and accelerated (Bottom) approximations of orbital motion in binary sources. We present the unlensed sources (solid lines) and the lensed images (dashed + dotted lines). For each source, there is a major and a minor image. The minor image is seen around the lens, and the major image is seen around the source. The green line is the flux-weighted average of lensed source positions, as observed on the sky. Note that $mag_{S,pri}$ = 16 and $mag_{S,sec}$ = 17. In both panels, the primary’s proper source motion $\boldsymbol{\mu}_{\boldsymbol{S},\sun}$ is [6 mas yr-1, 3 mas yr-1]; the secondary’s proper source motion relative to the primary source $\boldsymbol{\mu}_{\boldsymbol{S_{s}},\sun}$ is [9 mas yr-1, 7 mas yr-1]. The acceleration of the secondary source $\mathbb{a}_{\boldsymbol{S\mathrm{rel}},\sun}$ is [0.5 mas yr-2, -2 mas yr-2] in the lower panel. In both cases, the Einstein time $t_{E,\sun}$=269 days, the Einstein radius is 4.9 mas
, and $u_{0,\sun}$ =1.01. The lens mass is $10M_{\odot}$, and it is held stationary.
IV.2 Linear and Accelerated Orbital Approximations
@@
where $\boldsymbol{\mu}_{\boldsymbol{S},\sun}$ is the proper motion of the primary source, and $\boldsymbol{\mu}_{\boldsymbol{S_{s}},\sun}$ is the proper motion of the secondary source calculated as $\boldsymbol{\mu}_{\boldsymbol{S_{s}},\sun}=\boldsymbol{\mu}_{\boldsymbol{S},\sun}+\boldsymbol{\Delta\mu}_{\boldsymbol{S_{s}},\sun}$. Furthermore, we account for the parallactic motion $\pi_{S}\boldsymbol{P}(t,\alpha,\delta)$ where $\pi_{S}$ is the parallax amplitude $1/d_{S}$, and $\boldsymbol{P}(t,\alpha,\delta)$ is the actual parallax direction and fractional amplitude on the sky at time $t$ in the direction of the source-lens system $(\alpha,\delta)$, given by the difference of the Earth and Sun’s position, normalized by 1 AU. A more complete description of how parallax is defined in BAGLE can be found in J. Lu et al. (submitted).
@@
An example of the linear and accelerated orbital approximations is presented in Figure 3, showing the relative change in the projected Right Ascension ($\Delta\alpha^{*}$) and Declination ($\Delta\delta$) of the primary source (lensed and unlensed), secondary source (lensed and unlensed), and the lens. Here, $\Delta\alpha{{}^{*}}=\Delta\alpha\cos{\delta}$. The lensed images are further categorized into the major (“+”) image and minor (“-”) image. These figures can be generated using BAGLE with the code lines in Table 1. For this example, we use BSPL_PhotAstrom_noPar_LinOrbs_Param2, which is a model for linear orbits. One can follow the same steps below to generate plots for accelerated orbits.
@@
3# Define model parameters and pass them into a model instance below.
4# mag_base and b_sff are arrays or lists. All other parameters are floats.
5# raL and decL must be defined when calling a parallax model.
6
7bsplorbits = model.BSPL_PhotAstrom_noPar_LinOrbs_Param2(
@@
17
18# Define time array
19t = np.arange(t0 - 10*tE, t0 + 10*tE, 1)
20
21# Get resolved astrometry for unlensed source positions
22xS_unlensed = bsplorbits.get_resolved_source_astrometry_unlensed(t)
23srce_pos_primary = xS_unlensed[:, 0, :]
24srce_pos_secondary = xS_unlensed[:, 1, :]
25
26# Get lens astrometry
27lens = bsplorbits.get_lens_astrometry(t)
28
29# Get unresolved astrometry for lensed source positions
30xS_lensed = bsplorbits.get_astrometry_shift(t)
31lensed_pos_pri = xS_lensed[:, 0, :]
32lensed_pos_sec = xS_lensed[:, 1, :]
Table 1: Example BAGLEcode used to initialize the BSPL orbital model and compute unlensed and lensed source astrometry.
@@
IV.3 Full Keplerian Solutions
Figure 4: Source and lens trajectories for a simulated binary-source, point-lens microlensing event involving Keplerian orbits. We present the unlensed sources (solid lines) and the lensed images (dashed + dotted lines). The green line is the flux-weighted average of lensed source positions, as observed on the sky. Note that $mag_{S,pri}$ = 16 and $mag_{S,sec}$ = 17. This event has the following Keplerian elements: $\omega_{pri}=30\degree$, $\Omega_{sec}=10\degree$, $\textit{i}=90\degree$, $e=0.6$, $\textit{P}=1000$ days, $\aleph_{pri}=3$ mas and $\aleph_{sec}=8$ mas. The Einstein radius is 2.9 mas. We simulate this event over $t_{E}=208.47$ days. The lens (8 $M_{\odot}$) is held stationary.
All binary source models with Keplerian motion in BAGLE input the following Keplerian orbital parameters: $\omega_{pri}$ (set to 0 for circular orbits), i, $\Omega_{sec}$, e (for elliptical orbits), P, $t_{p}$, $\aleph_{pri}$, $\aleph_{sec}$. These models input an additional quantity $\boldsymbol{\mu}_{\boldsymbol{S,com},\sun}$, which is the proper motion of the source system’s center of mass in the observer’s frame of reference. $\boldsymbol{X}_{\boldsymbol{S_{com},0},\sun}$ is the position of the binary source’s center of mass at $t_{0,\sun}=t_{prim,0,\sun}$. While BAGLE inputs $t_{0,\sun}=t_{prim,0,\sun}$ as the time of closest approach between the primary source and the lens, and $u_{0,\sun}$ as the closest approach between the primary source and the lens, it can convert between different coordinate systems (e.g., to use the closest approach between the source system center of mass and the lens). Conversions for $t_{0,\sun}$ and $u_{0,\sun}$ are presented in Appendix C and D.
After accounting for the proper motion of the center of mass, the primary and secondary source trajectories (including parallactic motion) are given by: